OpenStudy (anonymous):

Three missiles are fired at an enemy arsenal. The probabilities the individual missile will hit the arsenal are 0.75, 0.85, and 0.9. Find the probability that at least two of the missiles hit the arsenal.

OpenStudy (perl):

Let X = # of missiles that hit the arsenal P( at least two missiles hit arsenal) = P( X > = 2 ) = P ( X = 2) + P (X=3)

OpenStudy (perl):

the probability of at least two missles hitting arsenal is equal to probability of exactly two missles hitting plus probability of exactly 3 missiles hitting

OpenStudy (perl):

m1 = first missile hits arsenal m2= second missile hits arsenal m3= third missile hits arsenal P(m1) = .75 P(m2) = .85 P(m3)=.9

OpenStudy (perl):

P(X=2) = P( m1&m2 or m1&m3 or m2&m3) P(X=3) = P (m1&m2&m3)

OpenStudy (perl):

= .75*.85 + .75*.9 + .85*.9 + (.75*.85*.9)

OpenStudy (anonymous):

thanks alooooooooooot

OpenStudy (anonymous):

A professor gives a true-false exam consisting of thirty T-F questions. The questions whose answers are “true” are randomly distributed among the thirty questions. The pro- fessor thinks that 3 of the class are serious, and have correctly mastered the material, and 4 that the probability of a correct answer on any question from such students is 75%. The remaining students will answer at random. She glances at two questions from a test picked haphazardly. Both questions are answered correctly. What is the probability that this is the test of a serious student?